I have two questions involving solving the area of a rectangle.
Question #1:
The area of the rectangle is 32a^3b^4 square units. Find the width of the rectangle.
The length of the rectangle is 4a^2b.
What is the width of the rectangle?
Question #2 correlates to question 1.
The length of a table is three times as long as its width. Suppose the width of the table is 5x+2. What is an algebraic expression for the length of the table?

Answer by Alan3354(69443)

Question #1:
The area of the rectangle is 32a^3b^4 square units. Find the width of the rectangle.
The length of the rectangle is 4a^2b.
What is the width of the rectangle?
---
Area = L*W
W = Area/L
===========================================
Question #2 correlates to question 1.
The length of a table is three times as long as its width. Suppose the width of the table is 5x+2. What is an algebraic expression for the length of the table?
3W = 3*(5x+2)
There's no correlation to Q #1.